WHAT IS THE AREA OF A RECTANGLE WITH A WIDTH OF 12 FEET AND A LENGTH OF 10 FEET?
WHAT IS THE PERIMETER OF A TRIANGLE WHERE THE SIDES MEASURE 6,8,AND 10 INCHES?
WHAT WOULD BE THE RADIUS OF A CIRCLE WITH A DIAMETER OF 8 INCHES?

Answer by bucky(2189)

WHAT IS THE AREA OF A RECTANGLE WITH A WIDTH OF 12 FEET AND A LENGTH OF 10 FEET?
.
The area of a rectangle is found by multiplying the length times the width. The length is
normally considered to be the longest side of the rectangle. Note here that this problem
calls the width 12 feet, but the length is 2 feet shorter than that.
.
Anyhow the formula for the Area of a rectangle is:
.
A = L*W
.
where A is the area in square units and L and W are the length and width (both in the same
units of measurement). So in this problem we call W = 12 feet and L = 10 feet so that the
area is:
.
A = L*W = 10 * 12 = 120 square feet.
.
WHAT IS THE PERIMETER OF A TRIANGLE WHERE THE SIDES MEASURE 6,8,AND 10 INCHES?
.
The perimeter is the distance around a triangle. It is found by adding the lengths of
the three sides.
In this problem the perimeter of the given triangle is 6 + 8 + 10 = 24 inches.
.
WHAT WOULD BE THE RADIUS OF A CIRCLE WITH A DIAMETER OF 8 INCHES?
.
The length of the radius is half the diameter. The diameter is a line that goes from one
point on the circle and through the center of the circle to a point on the opposite
side of the circle. The radius is a line that goes from a point on the circle just to the
center of the circle.
.
For this problem, since the diameter is 8 inches the radius (which is half of the diameter)
must be 4 inches.